# Existence of Minimal Surfaces in Infinite Volume Hyperbolic $3$-manifolds

### v1

@@ 5. Schottky Manifolds
6. Manifolds with Rank One Cusps
In [Cos21], the existence of closed minimal surfaces in hyperbolic 3-manifolds was established, excluding the case of rank-1 cusps (2.1). In this section, we study this case. We start by constructing special mean-convex or mean-concave barrier surfaces for each end using the shrinkwrapping technique [CG06]. One perhaps surprising finding is that the presence of rank-1 cusps can actually make it easier to produce finite area minimal surfaces (see Remark 6.13.)
6.1. Construction of Barrier Surfaces in the Ends
With the proof of Marden’s Conjecture  [CG06, Ago04], it is established that each end of a hyperbolic 3-manifold with finitely generated fundamental group has a product structure (Section 2). Before this proof, Calegari and Gabai had to rely on highly technical assumptions to construct shrinkwrapping surfaces [CG06, Theorem 1.10]. However, the resolution of Marden’s Conjecture has significantly simplified the construction of these surfaces [Cos21, Section 2.4]. Here, we give this simplified version. In particular, given two geodesics $\alpha_{1}$ and $\alpha_{2}$ in the end $\mathcal{E}\simeq\mathcal{T}\times[0,\infty)$, the collection of surfaces separating $\alpha_{1}$ and $\alpha_{2}$ forms a special homology class, and the area minimizing surface (possibly defective) within this class corresponds to a shrinkwrapping surface $\Sigma$.
Note that if the end is not cusped, the existence of shrinkwrapping surfaces follows from [Cos21]. Hence, here we only consider cusped ends. In what follows, we will use the following notation and setup:
Notation and Setup
Let $\mathcal{M}$ be a hyperbolic $3$-manifold with rank-1 cusps and let $\mathcal{E}$ be a cusped end. Let $\mathcal{M}^{0}$ be the manifold obtained from $\mathcal{M}$ by removing all rank-$1$ and rank-$2$ cusps. Let $\widehat{\mathcal{C}}_{\mathcal{M}}$ be the corresponding relative compact core of $\mathcal{M}$ (see Section 2.3), which is a compact core of $\mathcal{M}$ that intersects each toroidal component of $\partial\mathcal{M}^{0}$ along an entire torus and intersects each annular component of $\partial\mathcal{M}^{0}$ along a single incompressible annulus (See Figure 1). Let the open surface $\mathcal{T}\subset\partial\widehat{\mathcal{C}}_{\mathcal{M}}$ represent the cusped end $\mathcal{E}\cong\mathcal{T}\times[0,\infty)$. Let $\{\mathcal{U}_{i}\}_{i=1}^{m}$ be the collection of rank-$1$ cusps in $\mathcal{M}$ with ¯$\overline{\mathcal{E}}\cap\mathcal{U}_{i}\neq\emptyset$, and let $\widehat{\mathcal{U}}=\bigcup_{i=1}^{m}\mathcal{U}_{i}$. Let $\mathcal{A}_{i}=\partial\widehat{\mathcal{C}}_{\mathcal{M}}\cap\mathcal{U}_{i}$ denote the corresponding incompressible annuli. Fix a parametrization for each rank-$1$ cusp $\mathcal{U}_{i}\simeq\gamma_{i}\times(-\infty,\infty)\times[0,\infty)$ where each annulus $\mathcal{A}_{i}$ parametrized as $\mathcal{A}_{i}=\gamma_{i}\times(-1,1)\times\{0\}$. Similarly, reparametrize the end $\mathcal{E}\simeq\mathcal{T}\times[1,\infty)$ with notation $\mathcal{T}_{t}=\mathcal{T}\times\{t\}$ such that $\partial(\overline{\mathcal{T}}_{t})=\bigcup_{i=1}^{m}\gamma_{i}\times\{t\}\times\{0\}$ for $t\geq 1$. Notice that with this parametrization, for each $t\geq 1$, the surface $\widehat{\mathcal{T}}_{t}=\mathcal{T}_{t}\bigcup_{i=1}^{m}(\gamma_{i}\times\{t\}\times[0,\infty))$ is a cusped complete surface in $\mathcal{M}$. Furthermore, let $\widehat{\mathcal{U}}^{n}=\bigcup_{i=1}^{m}\gamma_{i}\times(-\infty,\infty)\times(n,\infty)$ and let $\mathcal{M}^{n}=\mathcal{M}-\widehat{\mathcal{U}}^{n}$ which is a manifold with boundary where the rank-1 cusps near $\mathcal{E}$ is trimmed with different depths.
Special Homology Class $\mathcal{Y}_{\Lambda}$: For a given pair of geodesics $\Lambda=\{\alpha_{1},\alpha_{2}\}$ in $\mathcal{E}\simeq\mathcal{T}\times[1,\infty)$, we define the special homology class $\mathcal{Y}_{\Lambda}$ as follows: Let $\mathcal{T}^{\prime}$ be a connected surface separating $\alpha_{1}$ and $\alpha_{2}$ in $\mathcal{E}$ with $\mathcal{T}^{\prime}\cap\widehat{\mathcal{U}}=\widehat{\mathcal{T}}_{1}\cap\widehat{\mathcal{U}}$. In other words, we modify $\widehat{\mathcal{T}}_{1}$ by pushing it in $\mathcal{M}^{0}$ so that it separates $\alpha_{1}$ and $\alpha_{2}$ in $\mathcal{M}^{0}$ without modifying its tails in the cusped region $\widehat{\mathcal{U}}$. With the notation above, define the special homology class $\mathcal{Y}_{\Lambda}$ as the space of complete surfaces $\{\mathcal{S}\}$ in $\mathcal{E}$ satisfying that
  • $\mathcal{S}$ is separating $\alpha_{1}$ and $\alpha_{2}$, i.e., $\mathcal{M}-\mathcal{S}=\mathcal{M}_{in}\cup\mathcal{M}_{out}$ two components with $\alpha_{1}\subset\overline{\mathcal{M}}_{in}$ and $\alpha_{2}\subset\overline{\mathcal{M}}_{out}$,
  • For any $n$, $\mathcal{S}\cap\mathcal{M}^{n}$ is in the same relative homology class of $\mathcal{T}^{\prime}\cap\mathcal{M}^{n}$ in $(\mathcal{M}^{n},\partial\mathcal{M}^{n})$, i.e., $[\mathcal{S}\cap\mathcal{M}^{n}]=[\mathcal{T}^{\prime}\cap\mathcal{M}^{n}]\in H_{2}(\mathcal{M}^{n},\partial\mathcal{M}^{n})$.
In other words, $\mathcal{S}\cup\mathcal{T}^{\prime}$ separates a compact region $\Omega_{n}$ in $\mathcal{M}^{n}$ with $\partial\Omega_{n}-(\mathcal{T}^{\prime}\cup\mathcal{S})\subset\partial\mathcal{M}^{n}$ and $int(\Omega_{n})\cap(\alpha_{1}\cup\alpha_{2})=\emptyset$. Notice that this condition automatically implies for any surface $\mathcal{S}\in\mathcal{Y}_{\Lambda}$, $\mathcal{S}$ separates $\mathcal{E}$ into two components with $\mathcal{E}-\mathcal{S}=\mathcal{E}_{in}\cup\mathcal{E}_{out}$ such that $\alpha_{1}\in\mathcal{E}_{in}$ and $\alpha_{2}\in\mathcal{E}_{out}$.
Definition 6.1 (Shrinkwrapping Surfaces in the Ends).
Let $\mathcal{M}$ be a hyperbolic $3$-manifold, and let $\mathcal{E}\simeq\mathcal{T}\times[0,\infty)$ be an end of $\mathcal{M}$. Let $\Lambda=\{\alpha_{1},\alpha_{2}\}$ be a pair of geodesics in $\mathcal{E}$. Let $\mathcal{Y}_{\Lambda}$ be the special homology class of surfaces separating $\alpha_{1}$ and $\alpha_{2}$ in $\mathcal{E}$ described above. We call an area minimizing surface $\Sigma_{\Lambda}$ for $\mathcal{Y}_{\Lambda}$ a shrinkwrapping surface of $\mathcal{E}$. Such a surface realizes the infimal area over all surfaces in $\mathcal{Y}_{\Lambda}$ and may be singular along $\Lambda$.
Remark 6.2 (Regularity of Shrinkwrapping Surfaces).
Calegari and Gabai showed that a shrinkwrapping surface $\Sigma$ is smooth and minimal everywhere except at $\Sigma\cap\Lambda$. They further studied the behavior of $\Sigma$ near non-smooth intersection $\Sigma\cap\Lambda$ by relating it to the thin obstacle problem and proved the mean-convexity of the surfaces on the sides of the geodesics [CG06, Lemma 1.31].
We use the shrinkwrapping surfaces to construct barrier surfaces at each end defined as follows.
Definition 6.3 (Mean-Convex and Concave Barrier Surfaces in the Ends).
Let $\mathcal{M}$ be a hyperbolic $3$-manifold, and let $\mathcal{E}\simeq\mathcal{T}\times[1,\infty)$ be an end of $\mathcal{M}$. Let $\Sigma$ be a surface in $\mathcal{E}$ separating the end into two components $\mathcal{E}_{in}$ and $\mathcal{E}_{out}$ where $\mathcal{E}_{in}$ has finite volume. If $\mathcal{E}_{in}$ is piecewise mean convex ($\mathbf{H}$ points towards $\mathcal{E}_{in}$ along $\Sigma$) with interior angle less than $\pi$, then we call $\Sigma$ a mean-convex barrier surface in $\mathcal{E}$. Similarly, if $\mathcal{E}_{in}$ is piecewise mean-concave ($\mathbf{H}$ points away from $\mathcal{E}_{in}$ along $\Sigma$) with exterior angle less than $\pi$, then we call $\Sigma$ a mean-concave barrier surface in $\mathcal{E}$.
To use these shrinkwrapping surfaces as barriers at the ends of a hyperbolic 3-manifold $\mathcal{M}$, the first author used a pair of sufficiently far simple geodesics $\alpha_{1}$ (closer to the compact core) and $\alpha_{2}$ (farther from the compact core), ensuring that the shrinkwrapping surface $\Sigma$ could intersect only one of them. If $\Sigma$ intersects only $\alpha_{1}$, he proved that $\Sigma$ is a mean-convex barrier surface, and if it intersects only $\alpha_{2}$, $\Sigma$ is a mean-concave barrier surface [Cos21, Theorem 3.1 (Step 1)]. Note that if $\Sigma$ does not intersect any of the geodesics, then it is automatically smooth and minimal everywhere.
In this construction, to make sure that our shrinkwrapping surfaces do not intersect both the curves $\alpha_{1}$ and $\alpha_{2}$, we need a bounded diameter result for such minimal surfaces. When dealing with a geometrically infinite end with bounded geometry (i.e., a positive injectivity radius bound), the results from [Cos21, Lemma 2.1, Lemma 2.2] and [CG06, Lemma 1.15] provide a uniform diameter bound for a minimal (or shrinkwrapping) surface of fixed genus or fixed homology class. However, in the case of a geometrically infinite end with unbounded geometry, the short geodesics prevent having an injectivity radius bound. In this case, the first author proved the following result, assuming there is no rank-1 cusp.
Lemma 6.4.
[Cos21, Lemma 2.3] [Pointwise Diameter Bound for Closed Minimal Surfaces] Let $\mathcal{M}$ be a hyperbolic $3$-manifold with no rank-$1$ cusp, and let $\mathcal{E}\simeq\mathcal{T}\times[0,\infty)$ be a geometrically infinite end with unbounded geometry. let $\mathcal{K}$ be a compact set in $\mathcal{E}$. Let $\mathcal{Y}_{\mathcal{K}}$ be the space of closed surfaces in $\mathcal{E}$ in the homology type of $\mathcal{T}$ and with nontrivial intersection with $\mathcal{K}$. Let $\Sigma$ be a surface in $\mathcal{Y}_{\mathcal{K}}$ with area equal to the infimal area over all surfaces in $\mathcal{Y}_{\mathcal{K}}$. Then, there exists $C_{\mathcal{K}}>0$ such that $d_{\mathcal{K}}(\Sigma)=\sup\{d(x,\mathcal{K})\mid x\in\Sigma\}<C_{\mathcal{K}}$.

### v2

@@ 5. Schottky Manifolds
6. Hyperbolic $3$-Manifolds with Rank One Cusps
In [Cos21], the existence of closed minimal surfaces in hyperbolic 3-manifolds was established, excluding the case of rank-1 cusps (2.1). In this section, we study this case.
In the following, relatively incompressible refers to the incompressibility of a map $f:(S,\partial S)\rightarrow(M,P)$ where $S$ is a surface with boundary and $P$ are the annuli corresponding to rank-1 cusps linking geometric ends together in the topological ends of a tame hyperbolic $3$-manifold.
Definition 6.1.
We say a geometric end $\mathcal{E}$ of a tame hyperbolic manifold $\mathcal{M}$ is essential after compressions if after performing some disk compressions (possibly none) produces a relatively incompressible surface that is not homotopic to any geometric end other than possibly $\mathcal{E}$.
As an example, if $\mathcal{M}$ is neither a product nor a compression body then any end is essential after compressions. Moreover, if $\mathcal{M}$ is not a product then any relatively incompressible end is essential after compressions. Theorem 1.5 is a consequence of the next theorem.
Theorem 6.2.
Let $\mathcal{M}=\mathbb{H}^{3}/\Gamma$ be an infinite volume complete hyperbolic $3$-manifold with rank-1 cusps that is homeomorphic to the interior of a compact manifold. If $\mathcal{M}$ contains a geometrically finite geometric end $\mathcal{E}$ that is essential after compressions, then $\mathcal{M}$ contains a finite area, complete, embedded, stable minimal surface $\mathcal{S}$.